Solve the equation: 49y-21√y+2=0
I have tried squaring the equation to remove the √ but I don't think this is right.
Could you help please?
Thanks.

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

Solve the equation: 49y - 21sqrt(y) + 2=0
I have tried squaring the equation to remove the √ but I don't think this is right.
Sub x for sqrt(y)
--> 49x^2 - 21x + 2 = 0
(7x - 1)*(7x - 2) = 0
Then sub back, and square to get y.

Answer by richwmiller(17219)

Think of y as x^2 and sqrt(y) as x
49x^2-21x+2=0
(7x-2)*(7x-1)
now sub back
(7 sqrt(y)-2) (7 sqrt(y)-1) = 0